# Covariant effective spacetimes of spherically symmetric electro-vacuum with a cosmological constant

### v1

@@ II.3 Reduced formulation
Combing the evolution equation (18) with the constraint equation $\mathcal{G}=G^{EM}=0$ leads to that $P_{\Gamma}$ is a constant [46]
@@
which is independent on both $t$ and $x$. Moreover, the constraint equation $H_{x}^{EM}=0$ is automatically satisfied due to $H^{EM}_{x}$ being proportional to $\mathcal{G}$. Notice that both the remaining constraints and the right-hand sides of the Hamilton’s equations are independent of $\Gamma$. Hence, we can choose the partial gauge fixing [48]
@@
00 0 $=\partial_{x}\Psi+N\frac{E^{2}}{\left(E^{1}\right)^{3/2}}Q.$ (46)
@@
$\left\{\bar{H}_{x}(N^{x}_{1}),\bar{H}_{x}(N^{x}_{2})\right\}$ =$=\left\{H_{x}^{GR}(N^{x}_{1}),H_{x}^{GR}(N^{x}_{2})\right\}$
@@
$\left\{\bar{H}(N_{1}),\bar{H}(N_{2})\right\}$ =$=\left\{H^{GR}(N_{1}),H^{GR}(N_{2})\right\}$
@@
In the first step in Eq. (50c), we notice that $H^{GR}$ does not involve the derivatives of $K_{I}$, while $H^{Q}$ is only a function of $E^{I}$ but not involves the derivatives of $E^{I}$. Thus, $H^{Q}$ does not contribute to the Poisson bracket between the smeared Hamiltonian constraints. It is worth to note that the constraint algebra (50) for the coupling model is precisely the same as the one in the vacuum case. Then, the reduced Hamiltonian
@@
determines the corresponding Hamilton’s equations
$\dot{K_{1}}$ :=$:=\left\{K_{1},\bar{\mathbb{H}}\right\}=\left\{K_{1},{\mathbb{H}}^{GR}\right\}-N\frac{3GE^{2}}{2\left(E^{1}\right)^{5/2}}Q^{2},$ (52)
$\dot{K_{2}}$ :=$:=\left\{K_{2},\bar{\mathbb{H}}\right\}=\left\{K_{2},{\mathbb{H}}^{GR}\right\}+N\frac{G}{2\left(E^{1}\right)^{3/2}}Q^{2},$ (53)
$\dot{E^{1}}$ :=$:=\left\{E^{1},\bar{\mathbb{H}}\right\}=\left\{E^{1},{\mathbb{H}}^{GR}\right\},$ (54)
$\dot{E^{2}}$ :=$:=\left\{E^{2},\bar{\mathbb{H}}\right\}=\left\{E^{2},{\mathbb{H}}^{GR}\right\}.$ (55)
@@ II.4 Classical solutions
In this subsection, by solving Eq. (46), the constraint equations $\bar{H}=0$ and $\bar{H}_{x}=0$ and the Hamilton’s equations (52)-(55) under different gauge fixings, we will obtain the corresponding solutions. Based on the solutions, it is more obvious to understand the covariance from the viewpoint of coordinate transformation.
@@
00 0 $=\dot{E^{1}}=\left\{E^{1},\bar{H}[N]+\bar{H}_{x}[N^{x}]\right\},$ (58)
00 0 $=\dot{E^{2}}=\left\{E^{2},\bar{H}[N]+\bar{H}_{x}[N^{x}]\right\}.$ (59)
@@
In what follows, we aim to obtain the explicit solutions in different coordinates. On one hand, we take the Schwarzschild gauge
@@
$f_{c}(x)=1-\frac{2GM}{x}+\frac{GQ^{2}}{x^{2}}-\frac{\Lambda}{3}x^{2}.$ (70)
@@
$N^{2}=f_{c}(x).$ (71)
Thus, we obtain the solution in Schwarzschild coordinates with the line element (20) being
${\rm d}s^{2}=-f_{c}(x){\rm d}t^{2}+f_{c}(x)^{-1}{\rm d}x^{2}+x^{2}{\rm d}\Omega^{2}.$ (72)
@@
${\rm d}s^{2}$ $=-{\rm d}\tau^{2}+\left[{\rm d}x\pm\sqrt{1-f_{c}(x)}\,{\rm d}\tau\right]^{2}+x^{2}{\rm d}\Omega^{2},$ (76)
@@
$\frac{{\rm d}Y}{{\rm d}x}=\frac{Q}{x}\frac{\sqrt{1-f_{c}(x)}}{f_{c}(x)}.$ (79)
Hence, two vector potentials obtained in different gauges differ only in the gauge term ${\rm d}Y(x)$, implying the covariance of EM field.
@@ III Effective theory
In the above section, we introduce how EM field is coupled to gravity within the Hamiltonian framework of the classical GR. Additionally, the general covariance of the classical model was examined with detailed calculations. Now let us turn to the effective model of EM field coupled to effective quantum gravity proposed by [41, 42], and show its covariance.
@@ III.1 Covariance of effective theory
In these works, the diffeomorphism constraint $H_{x}$ which generates the spatial diffeomorphism transformation takes its classical form, while the Hamiltonian constraint is modified into an effective one, denoted by $H_{\rm eff}$, with quantum gravity effects included. They are assumed to satisfy the algebraic relation
@@
where $\mu$, which depends only on the gravitational degrees of freedom, appears in Eq. (80c) as a result of quantum gravity effects. To match this modified constraint algebra, the classical metric is replaced with an effective metric $g_{\rho\sigma}^{(\mu)}$, given by the line element:
@@
With these definitions and assumptions, it is turned out in [41, 42] that the metric $g_{\rho\sigma}^{(\mu)}$ is covariant, meaning that it transforms as
@@
In the vacuum case, $H_{\rm eff}$ is just the effective Hamiltonian constraint of gravity, denoted by $H_{\rm eff}^{GR}$. With restricting the number of terms for consideration, it is derived from the above two conditions (i) and (ii) that $H_{\rm eff}^{GR}$ must be given by the following equations [41, 42]
@@
Here $s_{a}$ are spacetime scalars defined by
$s_{1}=E^{1},\qquad\ s_{2}=K_{2},\qquad s_{3}=\frac{K_{1}}{E^{2}},$ (86)
$s_{4}=\frac{\partial_{x}E^{1}}{E^{2}},\qquad s_{5}=\frac{1}{E^{2}}\partial_{x}\left(\frac{\partial_{x}E^{1}}{E^{2}}\right).$ (87)
It is easy to show that the classical Hamiltonian constraint (6) can be recovered if we set the free function $\mathcal{R}$ to be
@@
which is a solution to Eq. (85) with $\mu=1$. In this paper, to incorporate the quantum gravity effects, we will take
@@
where $\mathcal{X}$ denotes an arbitrary function of $\mu$.
@@
To achieve this, we introduce
@@
with an arbitrary function $\mathcal{Y}(\mu)$. Using Eq. (85), it can be verified through straightforward calculations that the algebra generated by the total constraints reads
@@
$\{\mathcal{G}[\beta],H_{x}[N]\}$ $=-\mathcal{G}[N^{x}\partial_{x}\beta],$ (93b)
@@
Hence, the constraints $H_{x}$ and $H_{\rm eff}$ form a subalgebra as proposed by Eq. (80). Now let us show that the covariance conditions (i) and (ii) are satisfied for $H_{\rm eff}$. At first, the derivatives of $K_{1}$ are not contained in $H_{\rm eff}$ as that for $H_{\rm eff}^{EM}$, ensuring that the condition (i) is satisfied. Moreover, a direct calculation shows that
@@
Now let us turn to the covariance of EM field in effective case. Due to the factor $\mu$ in the constraint algebra (93), it is easy to show following the procedure given in Sec. II.2 that the vector potential $A_{\rho}$ is no longer covariant, i.e., Eq. (27) does not hold any more. To fix the problem, as done for the metric, we will introduce the effective vector potential, denoted by $A_{\rho}^{(\mu)}$, to restore covariance of EM field. In other words, we will define $A_{\rho}^{(\mu)}$ such that
@@
To get the correct expression of $A_{\rho}^{(\mu)}$, we need to repeat the derivation shown in Sec. II.2, but with the new constraint algebra (93). In this case, Eqs. (32), (33) and (42) become
@@
$\delta\Gamma=$ $\mathcal{L}_{\alpha\mathfrak{N}+\beta}\Gamma+\left(N\frac{\sqrt{\mu E^{1}}}{E^{2}}\Gamma+\Psi\right)\partial_{x}\alpha-\Gamma\mathfrak{N}^{\rho}\partial_{\rho}\alpha+\partial_{x}\lambda,$
@@
where we have used the EOM (23) with $H$ replaced by $H_{\rm eff}$ in the second step, and the fact that $\mu$ is a spacetime scalar in the third step. Combining all these results, we obtain an explicit form of $A_{\rho}^{(\mu)}$ as
@@ III.2 Effective dynamics
Let us follow the procedure introduced in the Secs. II.3 and II.4 to solve the dynamics of the effective model, determined by $H_{\rm eff}^{GR}$ in Eq. (84) with $\mathcal{R}$ given by (90) and $H_{\rm eff}^{EM}$ in Eq. (92). In this case of $\mathcal{R}\neq 0$, $M_{\rm eff}$ is no longer a Dirac observable, while it is in the case of $\mathcal{R}=0$ for the vacuum model without a cosmological constant considered in [41, 42].
@@
00 0 $=\dot{E^{1}}=\left\{E^{1},\bar{H}_{\rm eff}[N]+\bar{H}_{x}[N^{x}]\right\},$ (104)
00 0 $=\dot{E^{2}}=\left\{E^{2},\bar{H}_{\rm eff}[N]+\bar{H}_{x}[N^{x}]\right\}.$
Under the areal gauge (56) and (57), substituting Eq. (102) into Eq. (104), we have
@@
Here, it is used that $\mu$ is a function of $s_{1}$ and $M_{\rm eff}$ as a consequence of (85b). Equation (105) can be further simplified as
@@
It should be emphasized that the functions $M_{{\rm eff},s_{2}}$, $\mathcal{X}_{M}$ and $\mathcal{Y}_{M}$ in Eqs. (105) and (107) need to be evaluated under the areal gauge (56) and (57).
Now let us consider the effective Hamiltonian constraint. In the areal gauge (56) and (57), $\bar{H}_{\rm eff}=0$ can be rewritten as
@@
Under the gauge condition (56) and (57), $M_{\rm eff}$, $\mathcal{X}(\mu)$ and $\mathcal{Y}(\mu)$ become functions of $E^{2}$ and $K_{2}$. In other words, Eq. (108) establishes a relation between $K_{2}$ and $E^{2}$. To solve this equation, a new gauge, such as the Schwarzschild gauge (67) or the PG gauge (73), is needed. Once such a gauge is imposed, Eq. (108) can be solved, and $N$ and $N^{x}$ can be determined using Eq. (107). With $\mu$ obtained from these solutions, the effective metric (81) then can be constructed.

### v3

@@ II.3 Reduced formulation
Combining the evolution equation (18) with the constraint equation $\mathcal{G}=G^{EM}=0$ leads to that $P_{\Gamma}$ is a constant [49]
@@
which is independent of both $t$ and $x$. Moreover, $H_{x}^{EM}$ vanishes due to $H^{EM}_{x}$ being proportional to $\mathcal{G}$. Notice that both the remaining constraints and the right-hand sides of Hamilton’s equations are independent of $\Gamma$. Hence, we can choose the partial gauge fixing [52]
@@
$0$ $=\partial_{x}\Psi+N\frac{E^{2}}{\left(E^{1}\right)^{3/2}}Q.$ (33)
@@
$\left\{\bar{H}_{x}(N^{x}_{1}),\bar{H}_{x}(N^{x}_{2})\right\}$ $=\left\{H_{x}^{GR}(N^{x}_{1}),H_{x}^{GR}(N^{x}_{2})\right\}$
@@
$\left\{\bar{H}(N_{1}),\bar{H}(N_{2})\right\}$ $=\left\{H^{GR}(N_{1}),H^{GR}(N_{2})\right\}$
@@
In the first step in Eq. (37c), we notice that $H^{GR}$ does not involve the derivatives of $K_{I}$, while $H^{Q}$ is only a function of $E^{I}$ but does not involve the derivatives of $E^{I}$. Thus, $H^{Q}$ does not contribute to the Poisson bracket between the smeared Hamiltonian constraints. It is worthwhile to note that the constraint algebra (37) for the coupling model is precisely the same as the one in the vacuum case. Then, the reduced Hamiltonian
@@
determines corresponding Hamilton’s equations
$\dot{K_{1}}$ $:=\left\{K_{1},\bar{\mathbb{H}}\right\}=\left\{K_{1},{\mathbb{H}}^{GR}\right\}-N\frac{3GE^{2}}{2\left(E^{1}\right)^{5/2}}Q^{2},$ (39)
$\dot{K_{2}}$ $:=\left\{K_{2},\bar{\mathbb{H}}\right\}=\left\{K_{2},{\mathbb{H}}^{GR}\right\}+N\frac{G}{2\left(E^{1}\right)^{3/2}}Q^{2},$ (40)
$\dot{E^{1}}$ $:=\left\{E^{1},\bar{\mathbb{H}}\right\}=\left\{E^{1},{\mathbb{H}}^{GR}\right\},$ (41)
$\dot{E^{2}}$ $:=\left\{E^{2},\bar{\mathbb{H}}\right\}=\left\{E^{2},{\mathbb{H}}^{GR}\right\}.$ (42)
@@ II.4 Classical solutions
In this subsection, by solving Eq. (33), the constraint equations $\bar{H}=0$ and $\bar{H}_{x}=0$ and Hamilton’s equations (39)–(42) under different gauge fixings, we will obtain the corresponding solutions. Based on the solutions, it is more obvious to understand the covariance from the viewpoint of coordinate transformation.
@@
$0$ $=\dot{E^{1}}=\left\{E^{1},\bar{H}[N]+\bar{H}_{x}[N^{x}]\right\},$ (45)
$0$ $=\dot{E^{2}}=\left\{E^{2},\bar{H}[N]+\bar{H}_{x}[N^{x}]\right\}.$ (46)
@@
In what follows, we aim to obtain the explicit solutions in different coordinates. On the one hand, we take the Schwarzschild gauge
@@
$f_{\rm cl}(x)=1-\frac{2GM}{x}+\frac{GQ^{2}}{x^{2}}-\frac{\Lambda}{3}x^{2}.$ (57)
@@
$N^{2}=f_{\rm cl}(x).$ (58)
Thus, we obtain the classical solution in Schwarzschild coordinates with the line element (20) being
${\rm d}s^{2}=-f_{\rm cl}(x){\rm d}t^{2}+f_{\rm cl}(x)^{-1}{\rm d}x^{2}+x^{2}{\rm d}\Omega^{2}.$ (59)
@@
${\rm d}s^{2}$ $=-{\rm d}\tau^{2}+\left[{\rm d}x\pm\sqrt{1-f_{\rm cl}(x)}\,{\rm d}\tau\right]^{2}+x^{2}{\rm d}\Omega^{2},$ (63)
@@
$\frac{{\rm d}Y}{{\rm d}x}=\frac{Q}{x}\frac{\sqrt{1-f_{\rm cl}(x)}}{f_{\rm cl}(x)}.$ (66)
Hence, two vector potentials obtained in different gauges differ only in the gauge term ${\rm d}Y(x)$, implying the covariance of the EM field.
@@ III Effective theory
In the above section, we introduce how the EM field is coupled to gravity within the Hamiltonian framework of the classical GR. Additionally, the general covariance of the classical model was examined with detailed calculations. Now let us turn to the effective model of the EM field coupled to the effective quantum gravity proposed by [42, 43], and show its covariance.
@@ III.1 Covariance of effective theory
In these works, the diffeomorphism constraint $H_{x}$ that generates the spatial diffeomorphism transformation takes its classical form, while the Hamiltonian constraint is modified into an effective one, denoted by $H_{\rm eff}$, with quantum gravity effects included. They are assumed to satisfy the algebraic relation
@@
where $\mu$, which depends only on the gravitational degrees of freedom, appears in Eq. (67c) as a result of quantum gravity effects. To match this modified constraint algebra, the classical metric is replaced with an effective metric $g_{\rho\sigma}^{(\mu)}$, given by the line element
@@
It should be mentioned that $\mu$ may change sign, as observed in some black hole models [42]. At first glance, such a sign change may seem problematic, as it would suggest that a spacetime region with signature (,,+,+)(-,-,+,+) must be smoothly joined to a standard Lorentzian region. However, as discussed in our previous work [42], this situation does not actually arise. The spacetime points where $\mu=0$ correspond to coordinate singularities. To extend the spacetime across $\mu=0$, one must adopt a different coordinate system that remains regular at these points. Using such coordinates, one will glue the region with $\mu>0$ to the original Lorentzian region in a manner that yields a spacetime symmetric with respect to the hypersurface $\mu=0$, without introducing any region where $\mu<0$ (see also Secs. III.4.4 and III.5.1 for concrete examples related to this issue). [^1] A similar scenario was previously encountered in certain quantum cosmological models (see, e.g., [54]), where the metric, at least in specific cases, takes the form ${\rm d}s^{2}=-f{\rm d}\eta^{2}+g\delta_{ab}{\rm d}x^{a}{\rm d}x^{b}.$ Since $f$ is not positive definite, this may suggest a signature change to $(+,+,+,+)$ at $f=0$. As discussed in the main text, such predictions require more careful and nuanced analysis due to the degeneracy of the metric at $f=0$, even though it remains unclear whether these points correspond to coordinate singularities.
With these definitions and assumptions, it turns out in [42, 43] that the metric $g_{\rho\sigma}^{(\mu)}$ is covariant, meaning that it transforms as
@@
In the vacuum case, $H_{\rm eff}$ is just the effective Hamiltonian constraint of gravity, denoted by $H_{\rm eff}^{GR}$. It is noted by [42, 43] that $H_{\rm eff}^{GR}$, as a scalar density, must take the form $E^{2}F$ where $F$ is a function of some basic scalars constructed from by $E^{I}$, $K_{I}$, and their derivatives. With restricting $F$ to depend on the same basic scalars $s_{a}$ ($a=1,\ldots,5$) as in the classical theory, namely,
$s_{1}=E^{1},\qquad\ s_{2}=K_{2},\qquad s_{3}=\frac{K_{1}}{E^{2}},$ (71)
$s_{4}=\frac{\partial_{x}E^{1}}{E^{2}},\qquad s_{5}=\frac{1}{E^{2}}\partial_{x}\left(\frac{\partial_{x}E^{1}}{E^{2}}\right),$
it is derived from the above two conditions (i) and (ii) that $H_{\rm eff}^{GR}$ must be given by the following equations [42, 43]:
@@
Equation (73b) implies
$\mu=\mu(s_{1},M_{\rm eff}).$ (74)
See [55] for more details on general solutions to Eq. (73). It is easy to show that the classical Hamiltonian constraint (6) can be recovered if we set the free function $\mathcal{R}$ to be
@@
which is a solution to Eq. (73) with $\mu=1$.
In this paper, from this point onward, to incorporate the quantum gravity effects, we will take
@@
where $\mathcal{X}$ denotes an arbitrary function of $\mu$. It should be emphasized that this choice of $\mathcal{R}$ in Eq. (77) is still a function of $s_{1}$ and $M_{\rm eff}$ by taking into account Eq. (74). Moreover, we allow $M_{\rm eff}$ to take any general form satisfying Eq. (3.6), and the analysis will proceed accordingly.
@@
To achieve this, we introduce the ansatz [^2] Here, this ansatz is introduced by modifying the classical expression (7) by a factor involving $\mu$. In particular, $\sqrt{\mu}$ is introduced in the second term to ensure that the Poisson bracket between $H_{\rm eff}$ and the Gauss constraint $\mathcal{G}$ retains the same structure as in the classical case, under the replacement $E^{1}\to\mu E^{1}$ [see Eq. (80c)]. The arbitrary function $\mathcal{Y}$ can be introduced because it will not change the structure of the constraint algebra (see Appendix B).
@@
with an arbitrary function $\mathcal{Y}(\mu)$. As shown in Appendix B, the algebra generated by the total constraints reads
@@
$\{\mathcal{G}[\beta],H_{x}[N^{x}]\}$ $=-\mathcal{G}[N^{x}\partial_{x}\beta],$ (80b)
@@
Here, it should be stressed that this algebra holds for all $H_{\rm eff}^{GR}$ and $H^{EM}_{\rm eff}$ with arbitrary selections of the free functions $\mathcal{X}$, $\mathcal{Y}$, and $M_{\rm eff}$ satisfying Eq. (73), according to the calculations given in Appendix B. According to Eq. (80), the constraints $H_{x}$ and $H_{\rm eff}$ form a subalgebra as proposed by Eq. (67). Now let us show that the covariance conditions (i) and (ii) are satisfied for $H_{\rm eff}$. At first, the derivatives of $K_{1}$ are not contained in $H_{\rm eff}$ as that for $H_{\rm eff}^{EM}$, ensuring that the condition (i) is satisfied. Moreover, a direct calculation shows that
@@
Now let us turn to the covariance of the EM field in the effective case. Due to the factor $\mu$ in the constraint algebra (80), it is easy to show following the procedure given in Sec. II.2 that the vector potential $A_{\rho}$ is no longer covariant, i.e., Eq. (25) does not hold any more. To fix the problem, as done for the metric, we will introduce the effective vector potential, denoted by $A_{\rho}^{(\mu)}$, to restore covariance of the EM field. In other words, we will define $A_{\rho}^{(\mu)}$ such that
@@
To get the correct expression of $A_{\rho}^{(\mu)}$, we need to repeat the derivation shown in Sec. II.2, but with the new constraint algebra (80). In this case, $\delta\Gamma$ in Eq. (28) and variation of the Lagrangian multipliers in Eq. (29) become
$\delta\Gamma=$ $\mathcal{L}_{\alpha\mathfrak{N}+\beta}\Gamma+\left(N\frac{\sqrt{\mu E^{1}}}{E^{2}}\Gamma+\Psi\right)\partial_{x}\alpha-\Gamma\mathfrak{N}^{\rho}\partial_{\rho}\alpha+\partial_{x}\lambda,$ (84)
$\delta N=$ $\mathcal{L}_{\alpha\mathfrak{N}+\beta}N+N\mathfrak{N}^{\rho}\partial_{\rho}\alpha,$
@@
where we have used the EOM (23) with $H$ replaced by $H_{\rm eff}$ in the second step and the fact that $\mu$ is a spacetime scalar in the third step. Combining all these results, we obtain an explicit form of $A_{\rho}^{(\mu)}$ as
@@ III.2 Effective dynamics
Let us follow the procedure introduced in Secs. II.3 and II.4 to solve the dynamics of the effective model, determined by $H_{\rm eff}^{GR}$ in Eq. (72) with $\mathcal{R}$ given by (77) and $H_{\rm eff}^{EM}$ in Eq. (79). In this case of $\mathcal{R}\neq 0$, $M_{\rm eff}$ is no longer a Dirac observable, while it is in the case of $\mathcal{R}=0$ for the vacuum model without a cosmological constant considered in [42, 43].
@@
$0$ $=\dot{E^{1}}=\left\{E^{1},\bar{H}_{\rm eff}[N]+\bar{H}_{x}[N^{x}]\right\},$ (91)
$0$ $=\dot{E^{2}}=\left\{E^{2},\bar{H}_{\rm eff}[N]+\bar{H}_{x}[N^{x}]\right\}.$
Under the conditions (43) and (44), substituting Eq. (89) into Eq. (91), we have
@@
Here, it is used that $\mu$ is a function of $s_{1}$ and $M_{\rm eff}$ as a consequence of Eq. (73b). Equation (92) can be further simplified as
@@
It should be emphasized that the functions $M_{{\rm eff},s_{2}}$, $\mathcal{X}_{M}$, and $\mathcal{Y}_{M}$ in Eqs. (92) and (94) need to be evaluated under the conditions (43) and (44).
Now let us consider the effective Hamiltonian constraint. In the conditions (43) and (44), $\bar{H}_{\rm eff}=0$ can be rewritten as
@@
Under the conditions (43) and (44), $M_{\rm eff}$, $\mathcal{X}(\mu)$, and $\mathcal{Y}(\mu)$ become functions of $E^{2}$ and $K_{2}$. In other words, Eq. (95) establishes a relation between $K_{2}$ and $E^{2}$. To solve this equation, a new gauge, such as the Schwarzschild gauge (54) or the PG gauge (60), is needed. Once such a gauge is imposed, Eq. (95) can be solved, and $N$ and $N^{x}$ can be determined using Eq. (94). With $\mu$ obtained from these solutions, the effective metric (68) then can be constructed.